Two groups of canoeists meet in the middle of a lake. After a brief visit, a person in canoe 1 pushes on canoe 2 to separate the canoes. Suppose the speeds of the two canoes after they are pushed apart are 0.60 m/s for canoe 1 and 0.46 m/s for canoe 2.

If the mass of canoe 1 is 310 kg , what is the mass of canoe 2?

____kg
Answer #1


let,


mass of the canoe, m1=310 kg


mass of the canoe is m2


speed of m1 is v1=0.6 m/sec


speed of m2 is v2=0.46 m/sec

by using law of conservation momentum,


m1*u1+m1*u2=m1*v1+m2*v2

0=310*0.6-m2*0.46


======> m2=404.35 kg


mass of the canoe, m2=404.35 kg